half the money collected at a show is donated to charity. Tickets cost $50 each. The charity received $3500. How many tickets were sold
Found 2 solutions by dkppathak, novice:
Answer by dkppathak(439)

half the money collected at a show is donated to charity. Tickets cost $50 each. The charity received $3500. How many tickets were sold
Let total number of ticket sold are X
cost of 1 ticket is $50
total amount collected at show =50X
amount deposited in charity will be 50x/2=25X
25x=3500
X=3500/25= 140
Answer total ticket sold at show are 140

Answer by novice(7)

very simple
charity*2=total turnover
$3500*$2=$7000
now divide $7000 by $50 =140tickets